# Classification of thin Jordan schemes

### v1

@@ 3.2.2. Relation to alternative algebras
A loop $(S,\diamond)$ is ring-alternative (an RA-loop, for short [4]) if the loop ring $RS$ is an alternative ring for any commutative associative ring $R$. It is known that each RA-loop is Moufang but the converse is not true [4]. The RA-loops were characterized in [4, Theorem 1] as follows.
@@
A non-associative loop $(S,\diamond)$ is an RA loop if and only if
  1. (1)
@@
Let $(S,\diamond)$ be a loop. The permutations $\ell_{a},a\in S$ form a thin Jordan scheme if and only if $(S,\diamond)$ is an RA loop.
@@
$\Leftarrow$” Let $(S,\diamond)$ be an RA loop. Then it is a Moufang loop [4] and, therefore, every element $a$ has an inverse $a^{-1}$ satisfying $\forall_{a,b\in S}\ a\diamond(a^{-1}\diamond b)=b$. This implies that $\ell_{a^{-1}}=\ell_{a}^{-1}=\ell_{a}^{t}$. Together with $\ell_{1_{\Omega}}=1_{S}$ we conclude that the partition $L:=\{\ell_{a}\}_{a\in S}$ of $S^{2}$ is a rainbow. If $\diamond$ is associative, then $(L,S)$ is a thin AS, and, therefore, is a Jordan scheme as well.
If $\diamond$ is non-associative, then being a Jordan scheme is equivalent to
@@
This is equivalent to the following relation in the loop ring $RS$:
@@
If $a,b,c$ associate in alphabetic order, then by Theorem 3.13 they associate in any order and (12) follows.
If $a,b,c$ do not associate, then by Theorem 3.13 $a\diamond(b\diamond c)=(b\diamond a)\diamond c,b\diamond(a\diamond c)=(a\diamond b)\diamond c$ and (12) holds again.
$\Rightarrow$” Assume now that the permutations $L=\{\ell_{u}\}_{u\in S}$ form a Jordan scheme. Picking $\omega_{0}=1_{\Omega}$ we obtain a binary operation $\odot$ on $L$ via
@@
Thus the mapping $a\mapsto\ell_{a}$ is an isomorphism between the magmas $(L,\odot)$ and $(S,\diamond)$ and for this reason we always have $\ell_{a}\odot\ell_{b}=\ell_{a\diamond b}$.
By Proposition 3.11 $(L,\odot)$ is a Moufang loop. We have to check the two conditions of Theorem 3.13. If $\ell_{u},\ell_{v},\ell_{w}$ associate, then they generate a subgroup222here we used the property of Moufang loop that subloop generated by three associate elements is associative of $L$, and, therefore, associate in any order. Thus the first requirement of Theorem 3.13 is satisfied.
To check the second condition of the Theorem consider a non-associate triple $\ell_{u},\ell_{v},\ell_{w}$, i.e. $\ell_{u}\odot(\ell_{v}\odot\ell_{w})\neq(\ell_{u}\odot\ell_{v})\odot\ell_{w}$. Then it follows from (8) that $\ell_{u}\odot(\ell_{v}\odot\ell_{w})=(\ell_{v}\odot\ell_{u})\odot\ell_{w}$ and $\ell_{v}\odot(\ell_{u}\odot\ell_{w})=(\ell_{u}\odot\ell_{v})\odot\ell_{w}$.
It remains to show that $(\ell_{u}\odot\ell_{v})\odot\ell_{w}=\ell_{u}\odot(\ell_{w}\odot\ell_{v})$. Since $\ell_{u},\ell_{v},\ell_{w}$ don’t associate, the elements $\ell_{w}^{-1},\ell_{v}^{-1},\ell_{u}^{-1}$ do not associate either. Now (8) implies $\ell_{w}^{-1}\odot(\ell_{v}^{-1}\odot\ell_{u}^{-1})=(\ell_{v}^{-1}\odot\ell_{w}^{-1})\odot\ell_{u}^{-1}$. Applying -1 to both sides we obtain $(\ell_{u}\odot\ell_{v})\odot\ell_{w}=\ell_{u}\odot(\ell_{w}\odot\ell_{v})$. Combining this with the equality proven in the previous paragraph we conclude that $(L,\odot)$ satisfies the second condition of the Theorem 3.13. Hence $(L,\odot)$ and, therefore, $(S,\diamond)$ are RA-loops. ∎
@@
The RA loops were classified in [9, Theorem 1.1]. To describe them we need to introduce a certain construction.
Let $G$ be a finite group satisfying $G/Z(G)\cong{\mathbb{Z}}_{2}^{2}$ and $G^{\prime}=\{e,s\}\subseteq Z(G)$ where $e$ is the identity element of $G$. Notice that according to [9] the condition $G/Z(G)\cong{\mathbb{Z}}_{2}^{2}$ always implies $G^{\prime}=\{e,s\}\subseteq Z(G)$.
Define an involution $g\mapsto g^{*},g\in G$ as follows
@@
On the set $G\times\{0,1\}$ we define a binary operation $\diamond$ as follows:
@@
where $g_{0}\in Z(G)$ is an arbitrary element chosen in advance. As in [9] we denote this algebraic structure as $L(G,*,g_{0})$.
Theorem 3.15.
([9, Theorem 1.1]) The algebraic structure $(L(G,*,g_{0}),\diamond)$ is a RA-loop and any RA-loop is isomorphic to $L(G,*,g_{0})$ for suitable $G$ and $g_{0}\in Z(G)$ as described earlier.
The statement below concentrates the properties of this loop which will be needed in what follows.
@@
    $Z(L)$ coincides with right and left nucleus of $L$
@@
Part (a). Write $g^{*}=g\alpha(g)$ where $\alpha(g)=e$ if $g\in Z(G)$ and $\alpha(g)=s$ otherwise. Then the property that $(gh)^{*}=h^{*}g^{*}$ is equivalent to $g,h]=\alpha(g)\alpha(h)\alpha(gh)^{-1}$.
It follows from the properties of $G$ that $g,h\in G$ do not commute if and only if the cosets $gZ(G),$ $hZ(G),$ $ghZ(G)$ are pairwise distinct. In this case $g,h]=s=\alpha(g)=\alpha(h)=\alpha(gh)$ and the required equality holds.
If $g,h$ commute, then they belong to an abelian normal subgroup $N:=\langle g,h,Z(G)\rangle\trianglelefteq G$. The restriction of $\alpha$ to $N$ is a homomorpism. Therefore $\alpha(gh)^{-1}\alpha(g)\alpha(h)=e=[g,h]$, hereby ending the proof of part (a).
Part (b). According to the product formula (13), for $g\in Z(G)$ and $h\in G$ we always have $gh=hg,g^{*}=g$. Therefore,
$(g,0)\diamond(h,y)=\left\{\begin{array}[]{cc}gh,&y=0\\ hg,&y=1\end{array}\right.=\left\{\begin{array}[]{cc}hg,&y=0\\ hg^{*},&y=1\\ \end{array}\right.=(h,y)\diamond(g,0)$
@@
It follows from this formula that $(g,0),g\in Z(G)$ commutes with any element of $L$. If $g\not\in Z(G)$, then $(g,0)$ and $(g,1)$ do not commute. Hence, neither $(g,0)$ nor $(g,1)$ is in $Z(L)$.
Part (c). Given that $L$ is an RA-loop, it is also a Moufang loop. Thus its left nucleus $N_{\lambda}(L)$ coincides with its right nucleus $N_{\rho}(L)$. Next we are going to show that $N_{\rho}(L)=Z(L)$. To this end, for $f,g,h\in G$, $x,y\in\{0,1\}$, we define
@@
If $f\in Z(G)$ then $f^{*}=f$ which together with $(hf)^{*}=f^{*}h^{*}$ implies $\mathsf{L}=\mathsf{R}$. Therefore, $Z(G)\times\{0\}\subseteq N_{\rho}(L)$.
If $f\not\in Z(G)$ then $f^{*}=fs\not\in Z(G)$ implying $hgf^{*}\neq hf^{*}g$ for any $g\in G$ not commuting with $f^{*}$. Therefore, $\mathsf{L}\neq\mathsf{R}$ for every $g$ with $g,f]\neq e$. Hence $(f,0)\not\in N_{\rho}(L)$ for any $f\not\in Z(G)$.
It remains to show that $(f,1)\not\in N_{\rho}(L)$ for any $f\in G$. Pick arbitrary $g,h\in G$ and compute $\mathsf{L}:=((g,1)\diamond(h,1))\diamond(f,1),\mathsf{R}:=(g,1)\diamond((h,1)\diamond(f,1))$. Applying (13) to computing $\diamond$-products we obtain
@@
It follows from $g_{0}\in Z(G)$ that $g_{0}^{*}=g_{0}$ and $\mathsf{L}=\mathsf{R}$ holds iff $fh^{*}g=gh^{*}f$. If $f\not\in Z(G)$, then choosing $h=e$ and $g\not\in C_{G}(f)$ we obtain $\mathsf{L}\neq\mathsf{R}$. If $f\in Z(G)$, then we can take $g,h\in G$ with $h^{*},g]\neq e$ to ensure $\mathsf{L}\neq\mathsf{R}$. Finally, we conclude that $(f,1)\not\in N_{\rho}(L)$ for any $f\in G$.
@@ 4. Coherent closure of thin Jordan schemes
The goal of this section is to show that the JS we’d found in the previous section are autonomous. For this purpose we compute the WL-closure $\mathfrak{C}$ of a Jordan scheme $\mathfrak{J}=(S,\{L_{s}\}_{s\in S})$ obtained from an RA-loop $(S,\diamond)$, compute the group of $\mathsf{JAut}(\mathfrak{C})$ and check that no algebraic fusion of $\mathfrak{C}$ coincides with $\mathfrak{J}$.
@@
Let $\Omega$ be a finite set and $G\leq{\sf Sym}(\Omega)$ a transitive subgroup. Then WL-closure of $G$ is a CC consisting of the $2$-orbits of the group $H:=C_{{\sf Sym}(\Omega)}(G)$.
@@
Let $\mathfrak{X}=(\Omega,C)$ be the coherent closure of $G$. Since $H$ centralizes $G$, each $h\in H$ is an automorphism of every $g\in G$. Therefore every basic relation of $\mathfrak{X}$ is a union of $2$-orbits of $H$. We have to show that each $2$-orbit of $H$ is a union of basic relations of $\mathfrak{X}$.
Pick an arbitrary pair $(\omega_{1},\omega_{2})\in\Omega^{2}$ and define $G_{\omega_{1}\rightarrow\omega_{2}}$ to be the set of all $g\in G$ which maps $\omega_{1}$ to $\omega_{2}$, that is $G_{\omega_{1}\rightarrow\omega_{2}}=\{g\in G\mid g(\omega_{1})=\omega_{2}\}=\{g\in G\mid(\omega_{1},\omega_{2})\in g\}$. Notice that $G_{\omega_{1}\rightarrow\omega_{2}}$ is non-empty and $G_{\omega_{1}\rightarrow\omega_{2}}=gG_{\omega_{1}}$ for any $g\in G_{\omega_{1}\rightarrow\omega_{2}}$.
The relation $r_{\omega_{1},\omega_{2}}:=\bigcap_{g\in G_{\omega_{1}\rightarrow\omega_{2}}}g$ is non-empty and is the union of certain elements from $C$. We claim that its is a $2$-orbit of $H$. Fix a permutation $g\in G_{\omega_{1}\rightarrow\omega_{2}}$ and write $r_{\omega_{1},\omega_{2}}=gr_{\omega_{1},\omega_{1}}$. Obviously $r_{\omega_{1},\omega_{1}}=1_{\Omega_{1}}$ where $\Omega_{1}=\mathsf{Fix}(G_{\omega_{1}})$. Clearly, $\Omega_{1}$ is $H$-invariant implying that $1_{\Omega_{1}}$ is $H$-invariant too. Notice that the action of $H$ on $1_{\Omega_{1}}$ is semi-regular (indeed, for any $\omega\in H$, if there existed $h\in H$ and $\omega^{\prime},\omega^{\prime\prime}\in H$ such that $h(\omega)=\omega$ and $h(\omega^{\prime})=\omega^{\prime\prime}$, then for any element $g\in G_{\omega\to\omega^{\prime}}$ we would have $g(\omega)=hgh^{-1}(\omega)=h(g(\omega))=\omega^{\prime\prime}$— a contradiction).
It is well-known that $|\Omega_{1}|=[N_{G}(G_{\omega_{1}}):G_{\omega_{1}}]=|C_{{\sf Sym}(\Omega)}(G)|=|H|$. Hence $H$ acts transitively in $1_{\Omega_{1}}$ and, therefore, it also acts transitively on $g1_{\Omega_{1}}=r_{\omega_{1},\omega_{2}}$, that is, $r_{\omega_{1},\omega_{2}}$ is a $2$ orbit of $H$. ∎
@@
Let $(S,\diamond)$ be a loop and $G\subseteq{\sf Sym}(S)$ the subgroup generated by $\{\ell_{s}\}_{s\in S}$. Then $C_{{\sf Sym}(S)}(G)$ coincides with the subgroup $\widehat{N}_{\rho}:=\{r_{s}\}_{s\in N_{\rho}}$ of ${\sf Sym}(S)$, where $N_{\rho}$ is the right nucleus of $(S,\diamond)$.
@@
Recall that $N_{\rho}:=\{z\in S\mid\forall_{x,y\in S}\ x\diamond(y\diamond z)=(x\diamond y)\diamond z\}$. It is easy to see that $N_{\rho}$ contains $1_{\Omega}$ and is $\diamond$-closed. Together with $r_{y\diamond z}=r_{y}r_{z}$, we obtain that $\widehat{N}_{\rho}$ is a subgroup of ${\sf Sym}(S)$. Notice that $\widehat{N}_{\rho}$ acts semiregularly on $S$.
A permutation $\phi\in{\sf Sym}(S)$ centralizes $G$ if and only if it commutes with $\ell_{s}$ for each $s\in S$, that is, $\phi(s\diamond u)=s\diamond\phi(u)$ for all $s,u\in S$. Substituting $u=1_{\Omega}$ to this equality, we conclude that $\phi=r_{\phi(1_{\Omega})}$, i.e. $\phi(x)=x\diamond\phi(1_{\Omega})$. Now the equality $\forall_{s,u\in S}\ \phi(s\diamond u)=s\diamond\phi(u)$ reads as $\forall_{s,u\in S}\ (s\diamond u)\diamond\phi(1_{\Omega})=s\diamond(u\diamond\phi(1_{\Omega}))$ implying that $\phi(1_{\Omega})\in N_{\rho}$. To sum up, $\Phi\in C_{{\sf Sym}(S)}(G)$ implies that $\phi=r_{\phi(1_{\Omega})}$ and $\phi(1_{\Omega})\in N_{\rho}$.

### v2

@@ 3.2.2. Relation to alternative algebras
A loop $(S,\diamond)$ is ring-alternative (an RA-loop, for short [4]) if the loop ring $RS$ is an alternative ring for any commutative associative ring $R$. It is known that every RA-loop is Moufang but the converse is not true [4]. RA-loops were characterized in [4, Theorem 1] as follows.
@@
A nonassociative loop $(S,\diamond)$ is an RA-loop if and only if the following hold:
  1. (a)
@@
Let $(S,\diamond)$ be a loop. The permutations $\{\ell_{a}\}_{a\in S}$ form a thin Jordan scheme if and only if $(S,\diamond)$ is an RA-loop.
@@
$\Leftarrow$” Let $(S,\diamond)$ be an RA-loop. Then $(S,\diamond)$ is a Moufang loop [4], and therefore every element $a$ has an inverse $a^{-1}$ satisfying $\forall_{a,b\in S}\ a\diamond(a^{-1}\diamond b)=b$. This implies that $\ell_{a^{-1}}=\ell_{a}^{-1}=\ell_{a}^{t}$. Together with $\ell_{1_{\Omega}}=1_{S}$, we conclude that the partition $L:=\{\ell_{a}\}_{a\in S}$ of $S^{2}$ is a rainbow. If $\diamond$ is associative, then $(L,S)$ is a thin association scheme, and therefore it is a Jordan scheme as well.
If $\diamond$ is nonassociative, then $(L,S)$ is a Jordan scheme if and only if
@@
Observe that this is equivalent to the following relation in the loop ring $RS$:
@@
If $a,b,c$ associate in lexicographic order, then by Theorem 3.13 they associate in any order and (11) follows. If $a,b,c$ do not associate, then by Theorem 3.13 $a\diamond(b\diamond c)=(b\diamond a)\diamond c,b\diamond(a\diamond c)=(a\diamond b)\diamond c$ and (11) holds once more.
$\Rightarrow$” Assume now that the permutations $L=\{\ell_{u}\}_{u\in S}$ form a Jordan scheme. Choosing $\omega_{0}=1_{\Omega}$ we obtain a binary operation $\odot$ on $L$ via
@@
Thus it always holds that $\ell_{a}\odot\ell_{b}=\ell_{a\diamond b}$. Therefore the mapping $a\mapsto\ell_{a}$ is an isomorphism between the magmas $(L,\odot)$ and $(S,\diamond)$.
By Proposition 3.11, $(L,\odot)$ is a Moufang loop. We must check the two conditions of Theorem 3.13. If $\ell_{u},\ell_{v},\ell_{w}$ associate, then they generate a subgroup222 Here we have used the fact that any subloop of a Moufang loop that is generated by three associate elements is associative. of $L$ and therefore associate in any order. Thus the first condition of Theorem 3.13 is satisfied.
To check the second condition, we consider a nonassociate triple $\ell_{u},\ell_{v},\ell_{w}$, i.e. $\ell_{u}\odot(\ell_{v}\odot\ell_{w})\neq(\ell_{u}\odot\ell_{v})\odot\ell_{w}$. Then it follows from (7) that $\ell_{u}\odot(\ell_{v}\odot\ell_{w})=(\ell_{v}\odot\ell_{u})\odot\ell_{w}$ and $\ell_{v}\odot(\ell_{u}\odot\ell_{w})=(\ell_{u}\odot\ell_{v})\odot\ell_{w}$.
It remains to show that $(\ell_{u}\odot\ell_{v})\odot\ell_{w}=\ell_{u}\odot(\ell_{w}\odot\ell_{v})$. Since $\ell_{u},\ell_{v},\ell_{w}$ do not associate, the elements $\ell_{w}^{-1},\ell_{v}^{-1},\ell_{u}^{-1}$ do not associate either. Now (7) implies $\ell_{w}^{-1}\odot(\ell_{v}^{-1}\odot\ell_{u}^{-1})=(\ell_{v}^{-1}\odot\ell_{w}^{-1})\odot\ell_{u}^{-1}$. Taking the inverse of both sides, we obtain $(\ell_{u}\odot\ell_{v})\odot\ell_{w}=\ell_{u}\odot(\ell_{w}\odot\ell_{v})$. Combining this with the equality of the previous paragraph, we conclude that $(L,\odot)$ satisfies the second condition of Theorem 3.13. Hence $(L,\odot)$, and therefore $(S,\diamond)$, are RA-loops. ∎
@@
RA-loops were classified in [9, Theorem 1.1]. To describe them, we must first introduce a certain construction.
Let $G$ be a finite group satisfying $G/Z(G)\cong{\mathbb{Z}}_{2}^{2}$ and $G^{\prime}=\{e,s\}\subseteq Z(G)$ where $e$ is the identity element of $G$. Note that according to [9], the property $G/Z(G)\cong{\mathbb{Z}}_{2}^{2}$ implies $G^{\prime}=\{e,s\}\subseteq Z(G)$.
Define an involution $g\mapsto g^{*}$ ($g\in G$) by
@@
We next fix an arbitrary element $g_{0}\in Z(G)$ and define a binary operation $\diamond$ on the set $G\times\{0,1\}$ as follows:
@@
As in [9], we denote this algebraic structure by $L(G,*,g_{0})$.
Theorem 3.15 ([9, Theorem 1.1]).
The algebraic structure $(L(G,*,g_{0}),\diamond)$ is a RA-loop. Moreover, any RA-loop is isomorphic to $L(G,*,g_{0})$ for suitable $G$ and $g_{0}\in Z(G)$ as described above.
The following result lists properties of $L(G,*,g_{0})$ that will be critical to our later arguments.
@@
    $Z(L)$ coincides with the right and left nucleus of $L$.
@@
Proof of (a). Define $g^{*}=g\alpha(g)$ where $\alpha(g)=e$ if $g\in Z(G)$ and $\alpha(g)=s$ otherwise. Then $(gh)^{*}=h^{*}g^{*}$ is equivalent to $g,h]=\alpha(g)\alpha(h)\alpha(gh)^{-1}$.
It follows from the properties of a group that $g,h\in G$ do not commute if and only if the cosets $gZ(G),$ $hZ(G),$ $ghZ(G)$ are pairwise distinct. In this case $g,h]=s=\alpha(g)=\alpha(h)=\alpha(gh)$, and the required equality holds.
On the other hand, if $g,h$ commute then they belong to an abelian subgroup $N:=\langle g,h,Z(G)\rangle$ that is normal in $G$. As the restriction of $\alpha$ to $N$ is a homomorpism, we have $\alpha(gh)^{-1}\alpha(g)\alpha(h)=e=[g,h]$, thereby completing the proof of part (a).
Proof of (b). For any $g\in Z(G)$ and $h\in G$, it follows from the product formula (12) that $gh=hg$, $g^{*}=g$. Therefore,
$(g,0)\diamond(h,y)=\left\{\begin{array}[]{cc}gh,&y=0\\ hg,&y=1\end{array}\right.=\left\{\begin{array}[]{cc}hg,&y=0\\ hg^{*},&y=1\\ \end{array}\right.=(h,y)\diamond(g,0),$
@@
From this it follows that $(g,0)$, $g\in Z(G)$, commutes with every element of $L$. If $g\not\in Z(G)$, then $(g,0)$ and $(g,1)$ do not commute. Hence, neither $(g,0)$ nor $(g,1)$ lies in $Z(L)$.
Proof of (c). Given that $L$ is an RA-loop, it is also a Moufang loop. Thus its left nucleus $N_{\lambda}(L)$ coincides with its right nucleus $N_{\rho}(L)$. We claim $N_{\rho}(L)=Z(L)$.
To this end, let $f,g,h\in G$, $x,y\in\{0,1\}$ and define
@@
If $f\in Z(G)$ then $f^{*}=f$, which together with $(hf)^{*}=f^{*}h^{*}$ implies $\mathsf{L}=\mathsf{R}$. Therefore, $Z(G)\times\{0\}\subseteq N_{\rho}(L)$.
If $f\not\in Z(G)$ then $f^{*}=fs\not\in Z(G)$, which implies $hgf^{*}\neq hf^{*}g$ for any $g\in G$ that does not commute with $f^{*}$. Therefore $\mathsf{L}\neq\mathsf{R}$ for every $g$ with $g,f]\neq e$. Hence $(f,0)\not\in N_{\rho}(L)$ for any $f\not\in Z(G)$.
It remains to show that $(f,1)\not\in N_{\rho}(L)$ for any $f\in G$. For arbitrary $g,h\in G$, we compute $\mathsf{L}:=((g,1)\diamond(h,1))\diamond(f,1),\mathsf{R}:=(g,1)\diamond((h,1)\diamond(f,1))$. Applying (12), we now obtain
@@
It follows from $g_{0}\in Z(G)$ that $g_{0}^{*}=g_{0}$. Thus $\mathsf{L}=\mathsf{R}$ holds iff $fh^{*}g=gh^{*}f$. If $f\not\in Z(G)$, then choosing $h=e$ and $g\not\in C_{G}(f)$ we obtain $\mathsf{L}\neq\mathsf{R}$. If $f\in Z(G)$, then we may take $g,h\in G$ with $h^{*},g]\neq e$ to ensure $\mathsf{L}\neq\mathsf{R}$. We thereby conclude that $(f,1)\not\in N_{\rho}(L)$ for any $f\in G$. ∎
@@ 4. Coherent closure of thin Jordan schemes
The goal of this section is to show that the Jordan schemes we have found in the previous section are autonomous. For this purpose we compute the WL-closure $\mathfrak{C}$ of the Jordan scheme $\mathfrak{J}=(S,\{\ell_{s}\}_{s\in S})$ obtained from an RA-loop $(S,\diamond)$. Following, we compute the group $\mathsf{JAut}(\mathfrak{C})$ and use it to check that no algebraic fusion of $\mathfrak{C}$ coincides with $\mathfrak{J}$.
@@
Let $\Omega$ be a finite set, and let $G\leq{\sf Sym}(\Omega)$ be a transitive subgroup. Then the WL-closure of $G$ is a coherent configuration consisting of the $2$-orbits of the group $H:=C_{{\sf Sym}(\Omega)}(G)$.
@@
Let $\mathfrak{X}=(\Omega,C)$ be the coherent closure of $G$. Since $H$ centralizes $G$, each $h\in H$ is an automorphism of every $g\in G$. Therefore every basic relation of $\mathfrak{X}$ is a union of $2$-orbits of $H$. We must show each $2$-orbit of $H$ is a union of basic relations of $\mathfrak{X}$.
Pick an arbitrary pair $(\omega_{1},\omega_{2})\in\Omega^{2}$ and define $G_{\omega_{1}\rightarrow\omega_{2}}$ to be the set of all $g\in G$ that map $\omega_{1}$ to $\omega_{2}$, that is $G_{\omega_{1}\rightarrow\omega_{2}}=\{g\in G\mid g(\omega_{1})=\omega_{2}\}=\{g\in G\mid(\omega_{1},\omega_{2})\in g\}$. Note that $G_{\omega_{1}\rightarrow\omega_{2}}$ is nonempty and $G_{\omega_{1}\rightarrow\omega_{2}}=gG_{\omega_{1}}$ for any $g\in G_{\omega_{1}\rightarrow\omega_{2}}$.
The relation $r_{\omega_{1},\omega_{2}}:=\bigcap_{g\in G_{\omega_{1}\rightarrow\omega_{2}}}g$ is nonempty and is the union of certain elements from $C$. We claim it is a $2$-orbit of $H$. Fix a permutation $g\in G_{\omega_{1}\rightarrow\omega_{2}}$ and write $r_{\omega_{1},\omega_{2}}=gr_{\omega_{1},\omega_{1}}$. Obviously $r_{\omega_{1},\omega_{1}}=1_{\Omega_{1}}$ where $\Omega_{1}=\mathsf{Fix}(G_{\omega_{1}})$. Clearly $\Omega_{1}$ is $H$-invariant implying that $1_{\Omega_{1}}$ is $H$-invariant as well.
Observe that the action of $H$ on $1_{\Omega_{1}}$ is semiregular. Otherwise, given any $\omega\in H$, if there were to exist $h,\omega^{\prime},\omega^{\prime\prime}\in H$ such that $h(\omega)=\omega$ and $h(\omega^{\prime})=\omega^{\prime\prime}$, then for any element $g\in G_{\omega\to\omega^{\prime}}$ we would obtain the contradiction $g(\omega)=hgh^{-1}(\omega)=h(g(\omega))=\omega^{\prime\prime}$.
Finally, it is well known that $|\Omega_{1}|=[N_{G}(G_{\omega_{1}}):G_{\omega_{1}}]=|C_{{\sf Sym}(\Omega)}(G)|=|H|$. Hence $H$ acts transitively in $1_{\Omega_{1}}$, and therefore it also acts transitively in $g1_{\Omega_{1}}=r_{\omega_{1},\omega_{2}}$. We conclude that $r_{\omega_{1},\omega_{2}}$ is a $2$ orbit of $H$. ∎
@@
Let $(S,\diamond)$ be a loop, and let $G\subseteq{\sf Sym}(S)$ be the subgroup generated by $\{\ell_{s}\}_{s\in S}$. Then $C_{{\sf Sym}(S)}(G)$ coincides with the subgroup $\widehat{N}_{\rho}:=\{r_{s}\}_{s\in N_{\rho}}$ of ${\sf Sym}(S)$, where $N_{\rho}$ is the right nucleus of $(S,\diamond)$.
@@
Recall that $N_{\rho}:=\{z\in S\mid\forall_{x,y\in S}\ x\diamond(y\diamond z)=(x\diamond y)\diamond z\}$. It is easy to see that $N_{\rho}$ contains $1_{\Omega}$ and is $\diamond$-closed. Together with $r_{y\diamond z}=r_{y}r_{z}$, we obtain that $\widehat{N}_{\rho}$ is a subgroup of ${\sf Sym}(S)$. Note that $\widehat{N}_{\rho}$ acts semiregularly on $S$.
A permutation $\phi\in{\sf Sym}(S)$ centralizes $G$ if and only if it commutes with $\ell_{s}$ for every $s\in S$, that is, if $\phi(s\diamond u)=s\diamond\phi(u)$ for all $s,u\in S$. Substituting $u=1_{\Omega}$ into this equality, we obtain $\phi=r_{\phi(1_{\Omega})}$, i.e. $\phi(x)=x\diamond\phi(1_{\Omega})$. Thus the equality $\forall_{s,u\in S}\ \phi(s\diamond u)=s\diamond\phi(u)$ becomes $\forall_{s,u\in S}\ (s\diamond u)\diamond\phi(1_{\Omega})=s\diamond(u\diamond\phi(1_{\Omega}))$ whence $\phi(1_{\Omega})\in N_{\rho}$. To summarize, $\Phi\in C_{{\sf Sym}(S)}(G)$ implies $\phi=r_{\phi(1_{\Omega})}$ and $\phi(1_{\Omega})\in N_{\rho}$.